# Lennart Berg and Torben Wiedemann

### v1

@@ 5 The action of Weyl elements in $H_{3}$-graded groups
Assume that $G=\operatorname{Ch}(D_{6},\mathcal{R})$ for a commutative ring $\mathcal{R}$ and that $(U_{\alpha})_{\alpha\in H_{3}}$ is the $H_{3}$-grading constructed in 4.10. Then a straightforward matrix computation (see 2.11) shows that for all $\alpha\in H_{3}$ and all $x,y\in\mathcal{R}$, the twisting involution is given by $\tilde{\theta}_{\alpha}^{H}(x,y)^{*}=\tilde{\theta}_{\alpha}^{H}(-x,y)$.
Remark 5.22.
Let $\xi,\zeta$ be two roots in $H_{4}$. Then either $\Psi_{\xi,\zeta}$ is of type $A_{1}$, $A_{2}$ or $A_{1}\times A_{1}$, or $\xi,\zeta$ lie in a common subsystem of type $H_{3}$. It follows from this observation that the twisting involution, defined in the same way as in 5.18, is also well-defined in $H_{4}$-graded groups, and that 5.20 holds in this case as well.
@@ 6 The parametrisation theorem
In this section, we fix a root base $\Delta=(\rho_{1},\rho_{2},\rho_{3})$ of $H_{3}$ in standard order and an $H_{3}$-graded group $(G,(U_{\alpha})_{\alpha\in H_{3}})$. For each $\delta\in\Delta$, we fix a $\delta$-Weyl element $w_{\delta}$ and we put $w_{-\delta}\coloneqq w_{\delta}^{-1}$. For any word $\bar{\delta}=(\delta_{1},\ldots,\delta_{k})$ over Δ($\Delta\cup(-\Delta)$, we put $w_{\bar{\delta}}\coloneqq w_{\delta_{1}}\cdots w_{\delta_{k}}$ and $\sigma_{\bar{\delta}}\coloneqq\sigma(\bar{\delta})\coloneqq\sigma_{\delta_{1}}\cdots\sigma_{\delta_{k}}$.
In this section, we state the parametrisation theorem for $\Phi$-graded groups for the special case $\Phi=H_{3}$ and give a sketch of its proof (6.15). More details and a complete proof of the parametrisation theorem in full generality can be found in [Wie23, Chapter 4]. Combining this theorem with the results from the previous section, we obtain a first parametrisation of $(G,(U_{\alpha})_{\alpha\in H_{3}})$, see 6.19. This parametrisation will be refined in the following sections using the blueprint technique.
@@
The crucial property of $H_{3}$ that simplifies the proof of the parametrisation theorem is that any pair $\alpha$, $\beta$ of orthogonal roots is automatically adjacent. In particular, any $\beta$-Weyl element centralises $U_{\alpha}$.
@@
If $\bar{\beta}$ is a $\Delta$-expression of a root $\beta\in H_{3}$, then $w_{\bar{\beta}}$ is a $\beta$-Weyl element by 3.9 (ii).
The purpose of the parametrisation theorem is to establish the existence of a parametrisation of $(G,(U_{\alpha})_{\alpha\in H_{3}})$. In order to make this precise, we need to introduce the notions of parity maps and parameter systems.
@@
Let $X$ be any set. For all $s,t\in X$ and $m\in\mathbb{N}_{0}$, we denote by $P_{m}(s,t)$ the word $(u_{1},\ldots,u_{m})$ with $u_{k}=s$ if $k$ is odd and $u_{k}=t$ if $k$ is even.
@@
Let $A$ be an abelian group. A $\Delta$-parity map with values in $A$ is a map $\eta\colon H_{3}\times\Delta\rightarrow A,(\alpha,\delta)\mapsto\eta_{\alpha,\delta}=\eta(\alpha,\delta)$. Given any $\Delta$-parity map $\eta$, we define $\eta(\alpha,-\delta)\coloneqq\eta(\alpha^{\sigma(\delta)},\delta)^{-1}$ for all $\alpha\in H_{3}$, $\delta\in\Delta$ and for any word $\bar{\delta}=(\delta_{1},\ldots,\delta_{m})$ over Δ($\Delta\cup(-\Delta)$, we define $\eta_{\alpha,\bar{\delta}}=\eta(\alpha,\bar{\delta})$ to be
$\prod_{i=1}^{m}\eta(\alpha^{\sigma(\delta_{1},\dots,\delta_{i-1})},\delta_{i})=\eta(\alpha,\delta_{1})\cdot\eta(\alpha^{\sigma(\delta_{1})},\delta_{2})\cdots\eta(\alpha^{\sigma(\delta_{1},\dots,\delta_{m-1})},\delta_{m})$
(with the convention that $\eta_{\alpha,\emptyset}=1_{A}$ for the empty word $\emptyset$). When $\Delta$ and $A$ are clear from the context, we will simply say that $\eta$ is a parity map. If $\eta(\alpha,P_{m}(\beta,\gamma))=\eta(\alpha,P_{m}(\gamma,\beta))$ for all $\alpha\in H_{3}$ and $\beta,\gamma\in\Delta$ where $m=m_{\alpha,\beta}$ denotes the group-theoretic order of $\sigma_{\beta}\sigma_{\gamma}$ in the Weyl group, then $\eta$ is called braid-invariant. If for all roots $\alpha,\beta\in H_{3}$ such that $\alpha$ is adjacent to $\beta$ and $-\beta$ and for all $\Delta$-expressions $\bar{\beta}$ of $\beta$, we have $\eta(\alpha,\bar{\beta})=1_{A}$, then $\eta$ is called adjacency-trivial. If for all $\alpha\in H_{3}$ and all $b\in A$, there exists a word $\bar{\delta}$ over $\Delta$ such that $\sigma_{\bar{\delta}}$ stabilises $\alpha$ and $\eta(\alpha,\bar{\delta})=b$, then $\eta$ is called complete.
@@
Since the Weyl group acts transitively on $H_{3}$, it suffices to verify completeness for one root $\alpha$.
@@
Let $(A,M)$ be a parameter system and let $\eta$ be a parity map with values in $A$. A parametrisation of $(G,(U_{\alpha})_{\alpha\in H_{3}})$ by $(A,M)$ with respect to $\eta$ and $(w_{\delta})_{\delta\in\Delta}$ is a family $(\theta_{\alpha}:M\rightarrow U_{\alpha})_{\alpha\in\Phi}$ of group isomorphisms such that for all $\alpha\in\Phi$, $\delta\in\Delta$ and $x\in M$, we have
@@
Let $A\coloneqq\{\pm 1\}^{2}$, let $\mathcal{R}$ be a commutative ring and put $M\coloneqq(\mathcal{R}^{2},+)$. Let $A$ act on $M$ via $(\varepsilon,\varepsilon^{\prime}).(x,y)\coloneqq(\varepsilon x,\varepsilon^{\prime}y)$ for all $x,y\in\mathcal{R}$ and $\varepsilon,\varepsilon^{\prime}\in\{\pm 1\}$. Then $(A,M)$ is a parameter system, called the standard parameter system for $\mathcal{R}$. It is faithful if and only if $2_{\mathcal{R}}\neq 0_{\mathcal{R}}$. We will show that every $H_{3}$-graded group is parametrised by a parameter system of this form.
The final missing ingredient in the parametrisation theorem is the notion of twisting groups.
@@
on the underlying set of $U_{\alpha}$ for each $\alpha\in H_{3}$. Then $A$ is called a twisting group for $(G,(w_{\delta})_{\delta\in\Delta})$ if for all $\alpha\in H_{3}$, $\delta\in\Delta\cup(-\Delta)$ and $b\in A$, we have $(b.x_{\alpha})^{w_{\delta}}=b.(x_{\alpha}^{w_{\delta}})$ for all $x_{\alpha}\in U_{\alpha}$. We will refer to the actions of $A$ on the root groups as twisting actions (on the root groups).
@@
Put $A\coloneqq\{\pm 1\}^{2}$. For all $\alpha\in H_{3}$ any $\alpha\in H_{3}$, we declare that $A$ acts on $U_{\alpha}$ via $(-1,-1).x_{\alpha}\coloneqq x_{\alpha}^{-1}$ and $(-1,1).x_{\alpha}\coloneqq x_{\alpha}^{*_{\alpha}}$ for all $x_{\alpha}\in U_{\alpha}$. Then by 5.19, $A$ is a twisting group for $(G,(w_{\delta})_{\delta\in\Delta})$. We will use this twisting group to construct a parametrisation of $G$. If $G=\operatorname{Ch}(D_{6},\mathcal{R})$ for some commutative ring $\mathcal{R}$ as in 4.10, then by 5.21, $(\varepsilon,\varepsilon^{\prime}).\tilde{\theta}_{\alpha}^{H}(x,y)=\tilde{\theta}_{\alpha}^{H}\big{(}(\varepsilon,\varepsilon^{\prime}).(x,y)\big{)}$ for all $x,y\in\mathcal{R}$ and all $(\varepsilon,\varepsilon^{\prime})\in A$.
@@
The following assertion is the main step in the proof of the parametrisation theorem.
Proposition 6.14 ([Wie23, 4.5.6]).
@@
Assume that $\eta$ is a complete braid-invariant $\Delta$-parity map with values in a twisting group $A$ for $(G,(w_{\delta})_{\delta\in\Delta})$ such that $G$ is square-compatible with respect to $\eta$. Let $\alpha\in H_{3}$ and let $\bar{\beta}$, $\bar{\gamma}$ be two words over Δ($\Delta\cup(-\Delta)$ such that $\alpha^{\sigma(\bar{\beta})}=\alpha^{\sigma(\bar{\gamma})}$. Put $b\coloneqq\eta_{\alpha,\bar{\beta}}\eta_{\alpha,\bar{\gamma}}^{-1}$. Then $x_{\alpha}^{w_{\bar{\beta}}}=(b.x_{\alpha})^{w_{\bar{\gamma}}}$ for all $x_{\alpha}\in U_{\alpha}$.

### v2

@@ 5 The action of Weyl elements in $H_{3}$-graded
Let $\mathcal{R}$ be a commutative ring, let $X\in\{D_{6},E_{8}\}$, put $\ell\coloneqq\operatorname{rank}(X)/2$, let $G=\operatorname{Ch}(X,\mathcal{R})$ be the Chevalley group from 4.22 or 4.23 and denote by $(\tilde{\theta}_{\alpha}^{H})_{\alpha\in H_{\ell}}$ the family of root isomorphisms from 4.26. Then a straightforward matrix computation (see 2.18) shows that for all $\alpha\in H_{\ell}$ and all $x,y\in\mathcal{R}$, the twisting involution is given by $\tilde{\theta}_{\alpha}^{H}(x,y)^{*}=\tilde{\theta}_{\alpha}^{H}(-x,y)$.
@@ 6 The parametrisation theorem
In this section, we fix a root system $\Phi$ with the property that any pair of orthogonal roots in $\Phi$ is adjacent and that the Weyl group acts transitively on $\Phi$. (In other words, $\Phi$ is irreducible of type $A$, $D$, $E$ or $H$.) We denote by $(G,(U_{\alpha})_{\alpha\in\Phi})$ a $\Phi$-graded group and by $\Delta$ a root base of $\Phi$ For each $\delta\in\Delta$, we fix a $\delta$-Weyl element $w_{\delta}$ and we put $w_{-\delta}\coloneqq w_{\delta}^{-1}$. For any word $\bar{\delta}=(\delta_{1},\ldots,\delta_{k})$ over Δ($\Delta\cup(-\Delta)$, we put $w_{\bar{\delta}}\coloneqq w_{\delta_{1}}\cdots w_{\delta_{k}}$ and $\sigma_{\bar{\delta}}\coloneqq\sigma(\bar{\delta})\coloneqq\sigma_{\delta_{1}}\cdots\sigma_{\delta_{k}}$.
The parametrisation theorem is a general tool which yields the existence of a parametrisation of $G$, that is, of a family of root isomorphisms which is “compatible” with the fixed family $(w_{\delta})_{\delta\in\Delta}$ of Weyl elements (see 6.11 for the precise definition). In this section, we state this theorem for $\Phi$-graded groups and give a self-contained proof (see 6.22). The restrictions on $\Phi$ in 6.1 simplify this task drastically. We will only be interested in the case $\Phi\in\{H_{3},H_{4}\}$, but there are no additional complications by allowing the other cases in 6.1. More details, motivation and a complete proof of the parametrisation theorem in full generality can be found in [Wie24, Chapter 4]. Combining this theorem with the results from the previous section, we obtain a parametrisation of $(G,(U_{\alpha})_{\alpha\in\Phi})$ for $\Phi\in\{H_{3},H_{4}\}$, see 6.24. This parametrisation will be refined to a standard coordinatisation (in the sense of 6.28) in the following sections using the blueprint technique.
@@
We will mostly use our assumption on orthogonal roots (see 6.1) in the following way: If $\alpha,\beta$ are orthogonal, then any $\beta$-Weyl element centralises $U_{\alpha}$.
@@
Let $\beta$ be a root. By the transitivity properties of the Weyl group, there exist $\Delta$-expressions of $\beta$. If $\bar{\beta}$ is a $\Delta$-expression of a root $\beta$, then $w_{\bar{\beta}}$ is a $\beta$-Weyl element by 3.10 (ii). In contrast, if $\bar{\beta}$ is any word over Δ($\Delta\cup(-\Delta)$ with $\sigma_{\bar{\beta}}=\sigma_{\beta}$, then $w_{\bar{\beta}}$ need not be a $\beta$-Weyl element because it is not clear that $w_{\bar{\beta}}\in U_{-\beta}U_{\beta}U_{-\beta}$. This complication is the reason why $\Delta$-expressions are more practical to work with.
In order to define parametrisations, we need to first introduce parity maps and parameter systems.
@@
Let $A$ be any set. For all $s,t\in A$ and $m\in\mathbb{N}_{0}$, we denote by $P_{m}(s,t)$ the word $(u_{1},\ldots,u_{m})$ with $u_{k}=s$ if $k$ is odd and $u_{k}=t$ if $k$ is even. We call $P_{m}(s,t)$ the braid word of $(s,t)$ of length $m$.
@@
Let $A$ be an abelian group. A $\Delta$-parity map with values in $A$ is a map $\eta\colon\Phi\times\Delta\rightarrow A,(\alpha,\delta)\mapsto\eta_{\alpha,\delta}=\eta(\alpha,\delta)$. Given any $\Delta$-parity map $\eta$, we define $\eta(\alpha,-\delta)\coloneqq\eta(\alpha^{\sigma(\delta)},\delta)^{-1}$ for all $\alpha\in\Phi$, $\delta\in\Delta$ and for any word $\bar{\delta}=(\delta_{1},\ldots,\delta_{m})$ over Δ($\Delta\cup(-\Delta)$, we define $\eta_{\alpha,\bar{\delta}}=\eta(\alpha,\bar{\delta})$ to be
$\prod_{i=1}^{m}\eta(\alpha^{\sigma(\delta_{1},\dots,\delta_{i-1})},\delta_{i})=\eta(\alpha,\delta_{1})\cdot\eta(\alpha^{\sigma(\delta_{1})},\delta_{2})\cdots\eta(\alpha^{\sigma(\delta_{1},\dots,\delta_{m-1})},\delta_{m}),$
with the convention that $\eta_{\alpha,\emptyset}=1_{A}$ for the empty word $\emptyset$. When $\Delta$ and $A$ are clear from the context, we will simply say that $\eta$ is a parity map. If $\eta(\alpha,P_{m}(\beta,\gamma))=\eta(\alpha,P_{m}(\gamma,\beta))$ for all $\alpha\in\Phi$ and $\beta,\gamma\in\Delta$ where $m=m_{\alpha,\beta}$ denotes the group-theoretic order of $\sigma_{\beta}\sigma_{\gamma}$ in the Weyl group, then $\eta$ is called braid-invariant. If for all roots $\alpha,\beta\in\Phi$ such that $\alpha$ is adjacent to $\beta$ and $-\beta$ and for all $\Delta$-expressions $\bar{\beta}$ of $\beta$, we have $\eta(\alpha,\bar{\beta})=1_{A}$, then $\eta$ is called adjacency-trivial. If for all $\alpha\in\Phi$ and all $b\in A$, there exists a word $\bar{\delta}$ over $\Delta$ such that $\sigma_{\bar{\delta}}$ stabilises $\alpha$ and $\eta(\alpha,\bar{\delta})=b$, then $\eta$ is called complete.
Example 6.7.
In 4.27, we have computed $\Delta$-parity maps for $\Phi=H_{3}$ and $\Phi=H_{4}$ with values in {±$\{\pm 1\}\times\{\pm 1\}$. We call them the standard parity maps for $H_{3}$ and $H_{4}$. The fact that they are “induced” by certain $\Phi$-graded groups ensures that they have many desirable properties (see 6.16). Still, there is nothing canonical about them: A twisting procedure as in 4.25 would allow us to obtain many more parity maps.
Note that the standard parity maps depend on the choice of a root base in $H_{3}$ or $H_{4}$. This is clear from the tables giving the values of the parity map, where each root $\alpha$ is expressed as a linear combination of a fixed root base in standard order. In the following, we will always tacitly assume that we choose the standard parity maps with respect to the root base $\Delta$ fixed in 6.1.
@@
By our assumption that the Weyl group acts transitively on $\Phi$, a parity map $\eta$ is complete if and only if for some $\alpha\in\Phi$ and all $b\in A$, there exists a word $\bar{\delta}$ over $\Delta$ such that $\sigma_{\bar{\delta}}$ stabilises $\alpha$ and $\eta(\alpha,\bar{\delta})=b$.
Remark 6.9.
Let $\alpha,\beta\in\Phi$. Then $\alpha$ is adjacent to $\beta$ and $-\beta$ if and only if $\alpha$ is orthogonal to $\beta$. Here the backwards implication holds by the assumptions in 6.1. This observation will be relevant for verifying that a parity map is adjacency-trivial.
@@
Let $(A,M)$ be a parameter system and let $\eta$ be a parity map with values in $A$. A parametrisation of $(G,(U_{\alpha})_{\alpha\in\Phi})$ by $(A,M)$ with respect to $\eta$ and $(w_{\delta})_{\delta\in\Delta}$ is a family $(\theta_{\alpha}:M\rightarrow U_{\alpha})_{\alpha\in\Phi}$ of group isomorphisms such that for all $\alpha\in\Phi$, $\delta\in\Delta$ and $x\in M$, we have
@@
Let $A\coloneqq\{\pm 1\}^{2}$, let $\mathcal{R}$ be a commutative ring and put $M\coloneqq(\mathcal{R}^{2},+)$. Let $A$ act on $M$ via $(\varepsilon,\varepsilon^{\prime}).(x,y)\coloneqq(\varepsilon x,\varepsilon^{\prime}y)$ for all $x,y\in\mathcal{R}$ and $\varepsilon,\varepsilon^{\prime}\in\{\pm 1\}$. Then $(A,M)$ is a parameter system, called the standard parameter system for $\mathcal{R}$. It is faithful if and only if $2_{\mathcal{R}}\neq 0_{\mathcal{R}}$. We will show in 9.33 that every root graded group of type $H_{3}$ or $H_{4}$ is parametrised by a parameter system of this form.
The final missing ingredient to formulate the parametrisation theorem is the notion of twisting groups.
@@
on the underlying set of $U_{\alpha}$ for each $\alpha\in\Phi$. Then $A$ is called a twisting group for $(G,(w_{\delta})_{\delta\in\Delta})$ if for all $\alpha\in\Phi$, $\delta\in\Delta\cup(-\Delta)$ and $b\in A$, we have $(b.x_{\alpha})^{w_{\delta}}=b.(x_{\alpha}^{w_{\delta}})$ for all $x_{\alpha}\in U_{\alpha}$. We will refer to the actions of $A$ on the root groups as twisting actions (on the root groups).
@@
Assume that $\Phi\in\{H_{3},H_{4}\}$ and put $A\coloneqq\{\pm 1\}\times\{\pm 1\}$. For all $\alpha\in\Phi$, we declare that $A$ acts on $U_{\alpha}$ via $(-1,-1).x_{\alpha}\coloneqq x_{\alpha}^{-1}$ and $(-1,1).x_{\alpha}\coloneqq x_{\alpha}^{*_{\alpha}}$ for all $x_{\alpha}\in U_{\alpha}$ where $*_{\alpha}$ denotes the twisting involution from 5.18 and 5.21. Then by 5.19 and 5.21, $A$ is a twisting group for $(G,(w_{\delta})_{\delta\in\Delta})$, called the standard twisting group for $(G,(w_{\delta})_{\delta\in\Delta})$. We will use this twisting group to construct a parametrisation of $G$.
Now let, in addition, $\mathcal{R}$ be a commutative ring, let $X\in\{D_{6},E_{8}\}$, put $\ell\coloneqq\operatorname{rank}(X)/2$, let $G=\operatorname{Ch}(X,\mathcal{R})$ be the Chevalley group from 4.22 or 4.23 and denote by $(\tilde{\theta}_{\alpha}^{H})_{\alpha\in H_{\ell}}$ the family of root isomorphisms from 4.26. Then by 5.22 and 5.21, $(\varepsilon,\varepsilon^{\prime}).\tilde{\theta}_{\alpha}^{H}(x,y)=\tilde{\theta}_{\alpha}^{H}\big{(}(\varepsilon,\varepsilon^{\prime}).(x,y)\big{)}$ for all $x,y\in\mathcal{R}$ and all $(\varepsilon,\varepsilon^{\prime})\in A$.
@@
We now show that the standard parity maps have all the desirable properties that we have defined in this section.
Proposition 6.16.
Let $X\in\{D_{6},E_{8}\}$, put $\ell\coloneqq\operatorname{rank}(X)/2$ and let $\eta\colon H_{\ell}\times\Delta\rightarrow\{\pm 1\}\times\{\pm 1\}$ be the standard parity map for $H_{\ell}$. Then the following hold:
  1. (a)
    Let $\mathcal{R}$ be a commutative ring, consider the Chevalley group $\operatorname{Ch}(X,\mathcal{R})$ with root groups $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{\ell}}$ from 4.22 or 4.23 and let $(w_{\delta}^{\prime})_{\delta\in\Delta}$ denote the family of Weyl elements from 4.27. Then the family $(\tilde{\theta}_{\alpha}^{H})_{\alpha\in H_{\ell}}$ of root isomorphisms from 4.26 is a parametrisation of $(\operatorname{Ch}(X,\mathcal{R}),(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{\ell}})$ by the standard parameter system for $\mathcal{R}$ (see 6.12) with respect to $\eta$ and $(w_{\delta}^{\prime})_{\delta\in\Delta}$.
  2. (b)
    $\eta$ is braid-invariant, adjacency-trivial, complete and satisfies $\eta(\alpha,\delta)=\eta(-\alpha,\delta)$ for all $\alpha\in H_{\ell}$, $\delta\in\Delta$.
  3. (c)
    Regard $\eta$ as having values in the standard twisting group $A=\{\pm 1\}\times\{\pm 1\}$ for $(G,(w_{\delta})_{\delta\in\Delta})$ from 6.14. Then $G$ is square-compatible with respect to $\eta$ and $(w_{\delta})_{\delta\in\Delta}$.
Proof.
Assertion a holds by (4.28), which is the defining property of $\eta$.
In the remaining part of the proof, we choose any commutative ring $\mathcal{R}$ with $2_{\mathcal{R}}\neq 0_{\mathcal{R}}$ and we fix the notation from a. To show that $\eta$ is adjacency-trivial, let $\alpha,\beta\in H_{\ell}$ such that $\alpha$ is adjacent to $\beta$ and $-\beta$ and let $\bar{\beta}$ be a $\Delta$-expression of $\beta$. Then $\alpha^{\sigma(\bar{\beta})}=\alpha^{\sigma(\beta)}=\alpha$ by 6.9 and $w_{\bar{\beta}}$ acts trivially on $U_{\alpha}$ by 6.2. Thus by a,
$\tilde{\theta}_{\alpha}^{H}(r,s)=\tilde{\theta}_{\alpha}^{H}(r,s)^{w_{\bar{\beta}}}=\tilde{\theta}_{\alpha}^{H}\big{(}\eta_{\alpha,\bar{\beta}}.(r,s)\big{)}$
for all $r,s\in\mathcal{R}$. Since the standard parameter system is faithful by 6.12 and our assumption that $2_{\mathcal{R}}\neq 0_{\mathcal{R}}$, it follows that $\eta_{\alpha,\bar{\beta}}=(1,1)$. Hence $\eta$ is adjacency-trivial. In a similar way, it follows from the braid relations (3.18) that $\eta$ is braid-invariant.
Observe that $\eta(\rho_{2},(\rho_{1},\rho_{1}))=(-1,-1)$ and $\eta(\rho_{2},(\rho_{3},\rho_{3}))=(-1,1)$ by Figure 4 and that these values generate $A$. Hence $\eta$ is complete by 6.8. Further, we have already observed in 4.27 that $\eta(\alpha,\delta)=\eta(-\alpha,\delta)$ for all $\alpha\in H_{\ell}$, $\delta\in\Delta$. This finishes the proof of b.
In addition to the action of $A$ on the root groups of $G$, we let $A$ act on $\mathcal{R}^{2}$ as in 6.12. Since the group $(\operatorname{Ch}(X,\mathcal{R}),(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{\ell}})$ is parametrised with respect to $\eta$, we have
$\tilde{\theta}_{\alpha}^{H}(x,y)^{w_{\delta}^{\prime}w_{\delta}^{\prime}}=\tilde{\theta}_{\alpha}^{H}\big{(}\eta(\alpha,(\delta,\delta)).(x,y)\big{)}$
for all $x,y\in\mathcal{R}$. At the same time, the action of $(w_{\delta}^{\prime})^{2}$ on $\tilde{\theta}_{\alpha}^{H}(x,y)$ is completely described by 5.20 and 5.22. Since the parameter system $(A,\mathcal{R}^{2})$ is faithful, we conclude: $\eta(\alpha,(\delta,\delta))=(1,1)$ if $\Psi_{\alpha,\delta}$ is of type $A_{1}$ or $A_{1}\times A_{1}$; $\eta(\alpha,(\delta,\delta))=(-1,-1)$ if $\Psi_{\alpha,\delta}$ is of type $A_{2}$; $\eta(\alpha,(\delta,\delta))=(-1,1)$ if $\alpha$ is in involution position with respect to $\delta$ and $\eta(\alpha,(\delta,\delta))=(1,-1)$ if $\alpha$ is in inverted involution position with respect to $\alpha$. Since 5.20 applies to $(G,(U_{\alpha})_{\alpha\in H_{\ell}})$ as well, the previous statement and the definition of the action of $A$ on $(U_{\alpha})_{\alpha\in H_{\ell}}$ imply that $G$ is square-compatible with respect to $\eta$. ∎
We now state and prove the parametrisation theorem. At first, we need to introduce some terminology which captures the transformations occurring in the solution of the word problem for Coxeter groups.
Definition 6.17.
Let $f$, $g$ be two words over $\Delta$.
  1. (a)
    We say that $f$ is elementary braid-homotopic to $g$ if there exist $\alpha,\beta\in\Delta$ and (possibly empty) words $f_{1},f_{2}$ over $\Delta$ such that $f=f_{1}P_{m}(\alpha,\beta)f_{2}$ and $g=f_{1}P_{m}(\beta,\alpha)f_{2}$ where $m$ denotes the group-theoretic order of $\sigma_{\alpha}\sigma_{\beta}$ in the Weyl group and $P_{m}$ is defined as in 6.5.
  2. (b)
    We say that $f$ is elementary square-homotopic to $g$ if there exist $\alpha\in\Delta$ and (possibly empty) words $f_{1},f_{2}$ over $\Delta$ such that $f=f_{1}(\alpha,\alpha)f_{2}$ and $g=f_{1}f_{2}$, or such that $f=f_{1}f_{2}$ and $g=f_{1}(\alpha,\alpha)f_{2}$. We say that $f$ and $g$ are square-homotopic if they lie in the equivalence relation generated by the relation “elementary braid-homotopic”.
  3. (c)
    We say that $f$ and $g$ are homotopic if they lie in the equivalence relation generated by the relation “braid-homotopic or square-homotopic”.
Lemma 6.18 ([Wie24, 4.5.3, 4.5.4]).
@@
Let $\eta$ be a braid-invariant $\Delta$-parity map with values in a twisting group $A$ for $(G,(w_{\delta})_{\delta\in\Delta})$ such that $G$ is square-compatible with respect to $\eta$. Let $\bar{\beta},\bar{\beta}^{\prime}$ be two words over Δ($\Delta\cup(-\Delta)$ such that $\sigma(\bar{\beta})=\sigma(\bar{\beta}^{\prime})$. Then $w_{\bar{\beta}}$ acts on $U_{\alpha}$ by $\eta(\alpha,\bar{\beta})$ if and only if $w_{\bar{\beta}^{\prime}}$ acts on $U_{\alpha}$ by $\eta(\alpha,\bar{\beta}^{\prime})$.
Proof.
Since $G$ is square-compatible with respect to $\eta$, we may replace every letter of $\bar{\beta}$ and $\bar{\beta}^{\prime}$ in $-\Delta$ by its negative. Hence we may assume that $\bar{\beta},\bar{\beta^{\prime}}$ are words over $\Delta$. Since $\sigma(\bar{\beta})=\sigma(\bar{\beta}^{\prime})$, the words $\bar{\beta}$ and $\bar{\beta^{\prime}}$ are then homotopic. If they are square-homotopic, then the assertion holds because $G$ is square-compatible with respect to $\eta$. If they are braid-homotopic, the assertion holds by the braid relations for Weyl elements (see 3.18) and because $\eta$ is braid-invariant. It follows that the assertion holds in this special case. See [Wie24, 4.5.3, 4.5.4] for a more detailed proof. ∎
Lemma 6.19 ([Wie24, 4.5.5]).
Let $\eta$ be an adjacency-trivial, braid-invariant $\Delta$-parity map with values in a twisting group $A$ for $(G,(w_{\delta})_{\delta\in\Delta})$ such that $G$ is square-compatible with respect to $\eta$. Let $\alpha\in\Phi$ and let $\bar{\beta}$ be a word over Δ($\Delta\cup(-\Delta)$ such that $\alpha^{\sigma(\bar{\beta})}=\alpha$. Then $x_{\alpha}^{w_{\bar{\beta}}}=\eta_{\alpha,\bar{\beta}}.x_{\alpha}$ for all $x_{\alpha}\in U_{\alpha}$.
Proof.
Since $\alpha^{\sigma(\bar{\beta})}=\alpha$, there exist $\rho_{1},\ldots,\rho_{n}\in\Phi\cap\alpha^{\perp}$ with $\sigma_{\bar{\beta}}=\sigma_{\rho_{1}}\cdots\sigma_{\rho_{n}}$. By 6.18, we may replace $\bar{\beta}$ by any word $\bar{\gamma}$ with $\sigma(\bar{\beta})=\sigma(\bar{\gamma})$, so we may assume that $\bar{\beta}=(\bar{\gamma}^{1},\ldots,\bar{\gamma}^{n})$ where $\bar{\gamma}^{1},\ldots,\bar{\gamma}^{n}$ are $\Delta$-expressions of $\rho_{1},\ldots,\rho_{n}$, respectively. For all $i\in\{1,\ldots,n\}$, $\eta(\alpha,\bar{\gamma}^{i})=1_{A}$ because $\eta$ is adjacency-trivial and further, $w_{\bar{\gamma}^{i}}$ centralises $U_{\alpha}$ by 6.4 and 6.2. Hence $w_{\bar{\beta}}$ centralises $U_{\alpha}$ and $\eta(\alpha,\bar{\beta})=1_{A}$, so the assertion holds. ∎
Note 6.20.
The roots $\rho_{1},\ldots,\rho_{n}$ in the proof of 6.19 need not be contained in $\Delta$, so we may not write $w_{\rho_{i}}$ and $\eta(\alpha,\rho_{i})$.
Proposition 6.21 ([Wie24, 4.5.6]).
Let $\eta$ be an adjacency-trivial, braid-invariant $\Delta$-parity map with values in a twisting group $A$ for $(G,(w_{\delta})_{\delta\in\Delta})$ such that $G$ is square-compatible with respect to $\eta$. Let $\alpha\in\Phi$ and let $\bar{\beta}$, $\bar{\gamma}$ be two words over Δ($\Delta\cup(-\Delta)$ such that $\alpha^{\sigma(\bar{\beta})}=\alpha^{\sigma(\bar{\gamma})}$. Put $b\coloneqq\eta_{\alpha,\bar{\beta}}\eta_{\alpha,\bar{\gamma}}^{-1}$. Then $x_{\alpha}^{w_{\bar{\beta}}}=(b.x_{\alpha})^{w_{\bar{\gamma}}}$ for all $x_{\alpha}\in U_{\alpha}$.
Proof.
This follows from 6.19. See [Wie24, 4.5.6] for details. ∎
The following result is precisely the content of the parametrisation theorem [Wie24, 4.5.16] for root systems $\Phi$ satisfying the restrictions in 6.1.
Theorem 6.22 ([Wie24, 4.5.16]).
Let $\eta$ be a complete, adjacency-trivial, braid-invariant $\Delta$-parity map with values in a twisting group $A$ for $(G,(w_{\delta})_{\delta\in\Delta})$ such that $G$ is square-compatible with respect to $\eta$. Then there exist an abelian group $(\mathcal{S},+)$, an action $\omega\colon A\times\mathcal{S}\rightarrow\mathcal{S},(b,r)\mapsto b.r$ of $A$ on the underlying set of $\mathcal{S}$ and a parametrisation $(\theta_{\alpha}\colon\mathcal{S}\rightarrow U_{\alpha})_{\alpha\in H_{3}}$ of $G$ by $(A,\mathcal{S})$ with respect to $\eta$ and $(w_{\delta})_{\delta\in\Delta}$ such that the twisting actions of $A$ on the root groups are induced by $\omega$. That is, we have $b.\theta_{\alpha}(s)=\theta_{\alpha}(b.s)$ for all $\alpha\in H_{3}$, $s\in\mathcal{S}$ and $b\in A$.